(2): A sample of helium at a pressure of 740. torr and in a volume of 2.58 L was heated from 24.0 to 750 °c. The volume of the container expanded to 2.81 L. What was the final pressure (in torr) of the helium? (3): What volume in liters does 0.136 g of O2 occupy at 20.0 °C and 2.36 atm? (4): A gas sample weighing 1.67 g has a volume of 276 mL at 589 torr and 28 °C. What is the molecular mass of the gas? (5): Propylene, CHs, reacts with hydrogen under pressure to give propane, CjHa How many liters of hydrogen (at 745 torr and 24 eC) react with 18.0 g of propylene?
Answer #1

(&X P, 740 tory 2 = 348 K 2- 3 48 29ד Prom sdaal gas eauahom 23 K mol. K 22 g/mol 589 .67 9 mol.K M = molar mass = 193

(5)

1 mole C3H6 = 42 g

1 mole C3H8 = 44 g

1 mole H2 = 2 g

From the balanced equation, 2g H2 is required to form 44 g C3H8.

Amount of H2 required to form 18 g C3H8 = 18 g C3H8 x 2 g H2 / 44 g C3H8 = 0.82 g

Now, from ideal gas equation,

PV = nRT

=> V = nRT/P = (0.82/2) x 0.0821 x 297 / (745/760)

=> V = 10.2 L